OpenStudy (anonymous):

length of a rectangle is 6yd longer than its width. If the perimeter of the rectangle is 48 yd, find its area

OpenStudy (anonymous):

w + 6 = l 2w +2l = 48 Solve simultaneously

OpenStudy (anonymous):

Wanna work on this together?

OpenStudy (anonymous):

l = 6 + w 2(6+w) + 2w = 48 sove to get l and w then area is simply l*w

OpenStudy (anonymous):

still dont get it

OpenStudy (anonymous):

Then let's work together, shall we? First of all, what letter do you want to use for our width?

OpenStudy (anonymous):

Because you know the length is 6 units longer than the width you know that: width + 6 = length

OpenStudy (anonymous):

or w+6 = l

OpenStudy (anonymous):

yep i get that

OpenStudy (anonymous):

Alright, now tell me what perimeter equals.

OpenStudy (anonymous):

ok, then you know that the perimeter of a rectangle is 2* width +2* length

OpenStudy (anonymous):

Like in general, what does perimeter mean?

OpenStudy (anonymous):

adding up all sides

OpenStudy (anonymous):

alright, so an equation for the perimeter (which we know is 48) would be....?

OpenStudy (anonymous):

Excellent. So L + L + W + W = perimeter, right? Or if we combine like terms, 2L + 2W = perimeter. Does that make sense?

OpenStudy (anonymous):

If it doesn't feel free to let me know. I can explain it more deeply.

OpenStudy (anonymous):

makes sense!

OpenStudy (anonymous):

Sweet! Ok, so in the beginning we decided that w+6 = L L is equal to (w+6) And now we want to say that 2L + 2W = perimeter. Do you know what L is equal to?

OpenStudy (anonymous):

(We don't know the exact number it equals, but give me a formula that tells us how big L is)

OpenStudy (anonymous):

-5?

OpenStudy (anonymous):

We haven't been told anything about that. Though, -5 would be a false value, as we can't have negative values. Hint: I actually gave you the answer in the same post where I asked you "Do you know what L is equal to?" Look carefully and you'll see what "L is equal to".

OpenStudy (anonymous):

w+6

OpenStudy (anonymous):

Excellent! So, here's the thing: Anytime you see L in another formula, you can completely replace the L with (w+6) So, what do we do here: 2L + 2W = perimeter

OpenStudy (anonymous):

(Hint: get rid of that L by replacing it)

OpenStudy (anonymous):

So, what did you get? I can't help if you don't reply. :P

OpenStudy (anonymous):

l = 6 + w 2(6+w) + 2w = 48 12+2w+2w=48 12+4w=48 -12 -12 4w=36 /4 /4 w=9 yd l=15 yd 9*15=135 yd^2 I am pretty sure that is your answer c:

OpenStudy (anonymous):

Width=x Length=x+6 Parameter=2(x+x+6)=48 ==>x=9 and x+6=15 Area=9*15=135

OpenStudy (anonymous):

"2(6+w) + 2w = 48" That is what I was hoping you would go for.